# Canonical Policy: Learning Canonical 3D Representation for Equivariant Policy

### v1

@@ Abstract
Visual Imitation learning has achieved remarkable progress in robotic manipulation, yet generalization to unseen objects, scene layouts, and camera viewpoints remains a key challenge. Recent advances address this by using 3D point clouds, which provide geometry-aware, appearance-invariant representations, and by incorporating equivariance into policy architectures to exploit spatial symmetries. However, existing equivariant approaches often lack interpretability and rigor due to unstructured integration of equivariant components. We introduce canonical policy, a principled framework for 3D equivariant imitation learning that unifies 3D point cloud observations under a canonical representation. We first establish a theory of 3D canonical representations, enabling equivariant observation-to-action mappings by grouping both in-distribution and out-of-distribution point clouds to a canonical representation. We then propose a flexible policy learning pipeline that leverages geometric symmetries from canonical representation and the expressiveness of modern generative models. We validate canonical policy on 12 diverse simulated tasks and 4 real-world manipulation tasks across 16 configurations, involving variations in object color, shape, camera viewpoint, and robot platform. Compared to state-of-the-art imitation learning policies, canonical policy achieves an average improvement of 18.0% in simulation and 37.6% in real-world experiments, demonstrating superior generalization capability and sample efficiency. For more details, please refer to the project website: https://zhangzhiyuanzhang.github.io/cp-website/.
Index Terms:
Equivariance, imitation learning, 3D vision.
@@ I Introduction
Imitation learning has made remarkable progress in robotic manipulation in recent years [1, 2, 3, 4, 5]. However, generalization and sample efficiency remain key challenges. In particular, visual imitation learning policies [1, 2] often struggle to generalize beyond the fixed dataset of human demonstrations on which they are trained. Unseen variations in object types, scene configurations, geometric layouts, or camera viewpoints can significantly degrade policy performance.
To improve generalization in visual imitation, recent research has explored the use of 3D point clouds as input observations [3, 6, 5, 7]. Point clouds encode the geometric structure of the environment and are invariant to visual distractors such as background textures or object appearances. As a result, policies trained on 3D representations can more effectively learn the mapping between geometry of observations and robot actions, thereby improving both generalization and sample efficiency.
Complementary to this, another line of work enhances generalization and sample efficiency to spatial configurations by embedding equivariance into the network architecture [8, 9, 10, 11]. Since robotic manipulation takes place in 3D Euclidean space and involves tasks that are often equivariant to rotations and translations, incorporating symmetry priors into the policy architecture can significantly improve both sample efficiency and generalization performance.
However, existing approaches that aim to achieve equivariant policy learning from 3D point clouds or 2D images often rely on re-designing the entire network using equivariant neural modules [8, 9, 10]. While this strategy introduces a certain degree of equivariance into the policy, it often lacks rigor, interpretability, and effectiveness in achieving equivariant learning. This is primarily because neural networks operate as black-box function approximators, and policy learning pipelines are typically long and complex, spanning from perception encoders to intermediate conditioning modules, and finally to generative action heads. This making it difficult to ensure that equivariance is preserved in a principled and verifiable manner throughout the entire pipeline. As a result, the geometric symmetry of the learned policy and its sample efficiency are often degraded.
@@
1. We propose a theory of 3D canonical representations for imitation learning. Through rigorous theoretical derivations, we demonstrate how both in-distribution and out-of-distribution point clouds can be efficiently grouped under a unified canonical distribution, and how this representation enables the construction of equivariant mappings from observations to actions. With this canonical representation, we show that 3D equivariant manipulation policies can be learned end-to-end from demonstrations, achieving strong spatial generalization in manipulation tasks.
@@
3. We propose a novel 3D encoder for policy learning that aggregates local neighborhood information during point cloud encoding. Experimental results show that it outperforms both MLP-based [3] and convolution-based 3D policy encoders [6, 7, 5]. This design improves the effectiveness of 3D policy learning.
4. We validate the effectiveness of canonical policy through extensive experiments. On a simulated benchmark of 12 diverse manipulation tasks, we compare against state-of-the-art 3D policy learning policies, including 3D diffusion policy (DP3)[3], improved 3D diffusion policy (iDP3)[6], and EquiBot [8]. Canonical policy consistently outperforms all baselines, achieving an average task success improvement of 18%. We further evaluate canonical policy on four real-world manipulation tasks across 16 configurations involving variations in object color, shape, camera viewpoint, and robot platform. Canonical policy consistently outperforms state-of-the-art methods, including diffusion policy (DP) [1], equivariant diffusion policy (EquiDiff) [9], and DP3 [3], achieving an average improvement of 37.6%.
@@ II-A Supervised Policy Learning
Recent advances in imitation learning have enabled robots to perform complex, long-horizon manipulation tasks using supervised training frameworks [1, 12, 2, 13]. This progress spans diverse areas, including scalable data collection pipelines [2, 13, 14], the exploration of expressive policy architectures such as diffusion [1], flow matching [15], and VQVAE [16], as well as large-scale models driven by vision-language foundations or expanded datasets [17, 18, 19].
Most existing work has leveraged visual modalities such as RGB images [1, 12, 2] or integrated tactile signals [20, 21, 22, 23, 24, 25, 26] to enhance physical interaction modeling. To improve robustness and generalization from limited demonstrations, recent work has explored structured data augmentation for visual policies, including semantic augmentations [27] and embodiment- or viewpoint-level transformations [28]. Meanwhile, 3D geometric representations such as point clouds [3, 6, 29, 7, 8, 30, 5] have shown particular promise for generalization. By capturing object shape while abstracting away distractors like color, background, and texture, 3D modalities allow policies to focus on task-relevant geometry. This enables more data-efficient learning and better transfer to unseen objects and environments, making 3D a particularly effective modality in scenarios that demand generalizable manipulation behavior.
@@ II-B Equivariance in Robot Learning
Robots operate in a three-dimensional Euclidean space, where manipulation tasks naturally exhibit geometric symmetries such as rotations. A growing body of work [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 9] demonstrates that incorporating symmetry priors into policy learning can significantly enhance sample efficiency and overall performance.
Moreover, point clouds are particularly well-suited as policy inputs for leveraging geometric symmetries, as they directly encode spatial structures and naturally transform under SE(3) operations. This alignment between input representation and task symmetry simplifies the design of equivariant policies and promotes better generalization across diverse object poses and environments. Recent works [10, 8, 11, 43] have explored leveraging the natural alignment between point cloud representations and geometric symmetries to design equivariant policies for improved sample efficiency and generalization.
@@ III Preliminaries
This section provides necessary background for our approach, which is situated in the context of visual-conditioned policy learning. We first describe the problem setup and formalize the input-output structure of the learning objective. We then introduce the concept of equivariance, which plays a central role in improving generalization and sample efficiency. Finally, we compare different visual modalities and discuss their implications for policy robustness and expressiveness.
@@ III-B Equivariance in Policy Learning
Equivariance is a desirable property in policy learning, especially for robotic manipulation tasks that involve geometric transformations. In manipulation, the same task (e.g., grasping, placing) often needs to be performed under different object poses or camera viewpoints. If the policy can inherently adapt to such variations through equivariance, it reduces the need for extensive data augmentation or retraining. Prior work [31, 33, 34, 8, 11] shows that incorporating symmetry priors significantly improves generalization and sample efficiency in robotic learning. Formally, a function $f$ is equivariant to a transformation $\mathbf{h}\in\mathrm{SE}(3)$ if
@@
III-C Different Visual Modalities
Visual representation plays a critical role in the perception module of policy learning, directly influencing the policy’s ability to generalize across varying task conditions. The choice of an appropriate visual modality is therefore essential for achieving robust and generalizable policy performance.
Image-based modalities, while widely used, are particularly sensitive to changes in appearance, such as lighting variations, object color, or texture, which can significantly degrade their performance under domain shifts.
In contrast, 3D representations, which explicitly encode geometric and positional information, offer greater robustness to such variations, enabling better appearance generalization across environments. For example, consider a robotic manipulation task where the goal is to identify and grasp a mug. When relying solely on 2D image representations, the model’s performance can degrade significantly under variations in lighting conditions, backgrounds, or object orientations, as these factors directly affect the object’s appearance. In contrast, 3D representations such as point clouds or meshes capture the object’s intrinsic geometric and spatial properties, which remain invariant to such changes. As a result, models utilizing 3D data can achieve more robust generalization across diverse environments.
Moreover, spatial transformations such as rotations and translations can be applied more naturally and accurately in 3D space than in 2D images. This property makes 3D representations especially suitable for tasks requiring geometric consistency and spatial reasoning.
In this work, we adopt point cloud as the input visual modality for policy learning. Point clouds are particularly well-suited for learning equivariant representations, as arbitrary transformations in the SE(3) group can be directly applied to the data without loss of geometric integrity.

### v2

# Canonical Policy: Learning Canonical 3D Representation for $\mathrm{SE}(3)$-Equivariant Policy

@@ Abstract
Visual Imitation learning has achieved remarkable progress in robotic manipulation, yet generalization to unseen objects, scene layouts, and camera viewpoints remains a key challenge. Recent advances address this by using 3D point clouds, which provide geometry-aware, appearance-invariant representations, and by incorporating equivariance into policy architectures to exploit spatial symmetries. However, existing equivariant approaches often lack interpretability and rigor due to unstructured integration of equivariant components. We introduce canonical policy, a principled framework for 3D equivariant imitation learning that unifies 3D point cloud observations under a canonical representation. We first establish a theory of 3D canonical representations, enabling equivariant observation-to-action mappings by grouping both seen and novel point clouds to a canonical representation. We then propose a flexible policy learning pipeline that leverages geometric symmetries from canonical representation and the expressiveness of modern generative models. We validate canonical policy on 12 diverse simulated tasks and 4 real-world manipulation tasks across 16 configurations, involving variations in object color, shape, camera viewpoint, and robot platform. Compared to state-of-the-art imitation learning policies, canonical policy achieves an average improvement of 18.0% in simulation and 39.7% in real-world experiments, demonstrating superior generalization capability and sample efficiency. For more details, please refer to the project website: https://zhangzhiyuanzhang.github.io/cp-website/.
@@ I Introduction
Imitation learning has made remarkable progress in robotic manipulation in recent years [diffusion_policy, zhao2023learning, DP3, o2024open, wang2024gendp]. However, generalization and sample efficiency remain key challenges. In particular, visual imitation learning policies [diffusion_policy, zhao2023learning] often struggle to generalize beyond the fixed dataset of human demonstrations on which they are trained. Unseen variations in object types, scene configurations, geometric layouts, or camera viewpoints can significantly degrade policy performance.
To improve generalization in visual imitation, recent research has explored the use of 3D point clouds as input observations [DP3, iDP3, wang2024gendp, huang20243d]. Point clouds encode the geometric structure of the environment and are invariant to visual distractors such as background textures or object appearances. As a result, policies trained on 3D representations can more effectively learn the mapping between geometry of observations and robot actions, thereby improving both generalization and sample efficiency.
Complementary to this, another line of work enhances generalization and sample efficiency to spatial configurations by embedding equivariance into the network architecture [Equibot, equidiff, EquiAct, gao2024riemann]. Since robotic manipulation takes place in 3D Euclidean space and involves tasks that are often equivariant to rotations and translations, incorporating symmetry priors into the policy architecture can significantly improve both sample efficiency and generalization performance.
However, existing approaches that aim to achieve equivariant policy learning from 3D point clouds or 2D images often rely on re-designing the entire network using equivariant neural modules [Equibot, equidiff, EquiAct]. While this strategy introduces a certain degree of equivariance into the policy, it often lacks rigor, interpretability, and effectiveness in achieving equivariant learning. This is primarily because neural networks operate as black-box function approximators, and policy learning pipelines are typically long and complex, spanning from perception encoders to intermediate conditioning modules, and finally to generative action heads. This makes it difficult to ensure that equivariance is preserved in a principled and verifiable manner throughout the entire pipeline. As a result, the geometric symmetry of the learned policy and its sample efficiency are often degraded.
@@
1. We propose a theory of 3D canonical representations for imitation learning. Through rigorous theoretical derivations, we demonstrate how both seen and novel point clouds can be efficiently grouped under a unified canonical representation, and how this representation enables the construction of equivariant mappings from observations to actions. With this canonical representation, we show that 3D equivariant manipulation policies can be learned end-to-end from demonstrations, achieving strong spatial generalization in manipulation tasks.
@@
3. We validate the effectiveness of canonical policy through extensive experiments. On a simulated benchmark of 12 diverse manipulation tasks, we compare against state-of-the-art 3D policy learning policies, including 3D diffusion policy (DP3)[DP3], improved 3D diffusion policy (iDP3)[iDP3], and EquiBot [Equibot]. Canonical policy consistently outperforms all baselines, achieving an average task success improvement of 18%. We further evaluate canonical policy on four real-world manipulation tasks across 16 configurations involving variations in object color, shape, camera viewpoint, and robot platform. Canonical policy consistently outperforms state-of-the-art methods, including diffusion policy (DP) [diffusion_policy], equivariant diffusion policy (EquiDiff) [equidiff], and DP3 [DP3], achieving an average improvement of 39.7%.
@@ II-A Supervised Policy Learning
Recent advances in imitation learning have enabled robots to perform complex, long-horizon manipulation tasks using supervised training frameworks [diffusion_policy, chi2024universal, zhao2023learning, fu2024mobile]. This progress spans diverse areas, including scalable data collection pipelines [zhao2023learning, fu2024mobile, wang2024dexcap], the exploration of expressive policy architectures such as diffusion [diffusion_policy], flow matching [flow_matching], and VQVAE [lee2024behavior], as well as large-scale models driven by vision-language foundations or expanded datasets [black2410pi0, team2023octo, zhao2024aloha].
Most existing work has leveraged visual modalities such as RGB images [diffusion_policy, chi2024universal, zhao2023learning] or integrated tactile signals [wang2024poco, yu2023mimictouch, xu2025unit, xue2025reactive, guzey2023see, lin2024learning, bhirangi2024anyskin] to enhance physical interaction modeling. To improve robustness and generalization from limited demonstrations, recent work has explored structured data augmentation for visual policies, including semantic augmentations [bharadhwaj2023roboagent] and embodiment- or viewpoint-level transformations [chen2024roviaug]. Meanwhile, 3D geometric representations such as point clouds [DP3, iDP3, xue2025demogen, huang20243d, Equibot, yang2025fp33dfoundationpolicy, wang2024gendp] have shown particular promise for generalization. By capturing object shape while abstracting away distractors like color, background, and texture, 3D modalities allow policies to focus on task-relevant geometry. This enables more data-efficient learning and better transfer to unseen objects and environments, making 3D a particularly effective modality in scenarios that demand generalizable manipulation behavior.
@@ II-B Equivariance in Robot Learning
Robots operate in a three-dimensional Euclidean space, where manipulation tasks naturally exhibit geometric symmetries such as rotations. A growing body of work [simeonov2023se, wang22onrobot, huang2023leveraging, pan2023tax, liu2023continual, jia2023seil, kim2023se, kohler2023symmetric, nguyen2024symmetry, eisner2024deep, hu2025push, zhao2025hierarchical, equidiff, actionflow, ipa] demonstrates that incorporating symmetry priors into policy learning can significantly enhance sample efficiency and overall performance.
Moreover, point clouds are particularly well-suited as policy inputs for leveraging geometric symmetries, as they directly encode spatial structures and naturally transform under SE(3) operations. This alignment between input representation and task symmetry simplifies the design of equivariant policies and promotes better generalization across diverse object poses and environments. Recent works [EquiAct, Equibot, gao2024riemann, huang2024matchpolicysimplepipeline] have explored leveraging the natural alignment between point cloud representations and geometric symmetries to design equivariant policies for improved sample efficiency and generalization.
@@ III Preliminaries
This section provides the necessary background for our approach, which is situated in the context of visual-conditioned policy learning. We first specify the problem setup and formalize the input–output structure of the learning objective. We then introduce equivariance, which plays a central role in improving generalization and sample efficiency. Next, we establish the theoretical foundation of our canonical policy by introducing canonical representations, which map all samples within the same equivariant group to a unified representation. Finally, we contrast canonicalization with $\mathrm{SE}(3)$ data augmentation and clarify the generalization benefits of canonical representations.
@@ III-B Equivariance in Policy Learning
Equivariance is a desirable property in policy learning, especially for robotic manipulation tasks that involve geometric transformations. In manipulation, the same task (e.g., grasping, placing) often needs to be performed under different object poses or camera viewpoints. If the policy can inherently adapt to such variations through equivariance, it reduces the need for extensive data augmentation or retraining. Prior work [simeonov2023se, huang2023leveraging, pan2023tax, Equibot, gao2024riemann] shows that incorporating symmetry priors significantly improves generalization and sample efficiency in robotic learning. Formally, a function $f$ is equivariant to a transformation $\mathbf{h}\in\mathrm{SE}(3)$ if
@@
III-C Theory of Canonical Representations
Figure 1: Illustraion of two distinct equivariant groups $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. Samples within the same group can be aligned via an $\mathrm{SE}$(3) transformation, while samples from different groups cannot be made to coincide.
Figure 2: Comparison of $\mathrm{SE}(3)$ augmentation and canonicalization: (a) $\mathrm{SE}(3)$ data augmentation; (b) $\mathrm{SE}(3)$ data canonicalization
Assumption 1.
A set of 3D point clouds can be partitioned into a finite number of mutually exclusive subsets: $\mathcal{D}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots,\mathcal{G}_{p}\},$ where $p$ is the total number of groups, $\mathcal{G}_{i}=\{\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},\cdots\mathbf{x}_{i}^{\ell}\}$ denotes the $i$-th group, and $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{3\times N}$ is the $\ell$-th point cloud sample in the group $\mathcal{G}_{i}$, consisting of $N$ points. For any two distinct groups $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$, it holds that
$\forall\mathbf{x}_{i}^{\ell}\in\mathcal{G}_{i},\;\mathbf{x}_{j}^{q}\in\mathcal{G}_{j},\;\forall\mathbf{R}\in\mathrm{SO}(3),\ \mathbf{b}\in\mathbb{R}^{3\times 1},\mathbf{1}\in\mathbb{R}^{1\times N}$
$\mathbf{x}_{i}^{\ell}\neq\mathbf{R}\mathbf{x}_{j}^{q}+\mathbf{b}\mathbf{1},$ (1)
where $\mathbf{R}$, $\mathbf{b}$, and $\mathbf{1}$ denote rotation matrix, translation vector, and all-ones vector, respectively. For notational simplicity, we use $\mathbf{x},\,\mathbf{y}$ to denote generic point cloud samples and omit group and sample indices. We also write $\mathbf{b}\mathbf{1}$ simply as $\mathbf{b}$ to represent the global translation. Moreover, within each group, we have
The illustration of the groups $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ is shown in Fig. 1. For any point cloud samples within the same group, there always exists an $\mathrm{SE}(3)$ transformation that can map one to the other. However, for point cloud samples belonging to two distinct groups $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$, no $\mathrm{SE}(3)$ transformations can make them coincide.
Remark 1.
Under Assumption 1, we can see that $\mathcal{G}_{i}$ forms an equivariant group, such that for any $\mathrm{SE}$(3)-equivariant function $f$, the following holds:
In the context of visual-conditioned policy learning, it is reasonable to assume that new observations from the same task can be related to samples in $\mathcal{D}$ through $\mathrm{SE}$(3) transformations, which is summarized below.
Assumption 2.
For a set of point clouds $\mathcal{D}$ collected from a demonstrated task, any point cloud observation from the same task but outside $\mathcal{D}$ can be mapped to a point cloud within $\mathcal{D}$ through an $\mathrm{SE}$(3) transformation.
In practice, the collected dataset $\mathcal{D}$ is typically diverse enough to cover a wide range of variations. Therefore, unseen point cloud observation but from the same task can be approximately mapped to $\mathcal{D}$ by $\mathrm{SE}$(3) transformations. For example, in robotic manipulation tasks, a dataset may include multiple demonstrations of grasping the same object from different orientations and viewpoints. These demonstrations inherently capture variations such as rotations, translations, and perspective shifts, all of which can be effectively modeled as rigid body transformations in 3D space. Consequently, unseen geometric layouts can often be mapped to known layouts in the dataset through appropriate $\mathrm{SE}$(3) transformations.
To simulate observations with new poses not present in the original dataset, we define a transformed variant using $\mathrm{SE}$(3) transformations. Specifically, for each group $\mathcal{G}_{i}$ and each sample $k$, let $\mathbf{R}_{i}^{k}\!\in\!\mathrm{SO}(3)$ and $\mathbf{b}_{i}^{k}\!\in\!\mathbb{R}^{3\times 1}$ be randomly sampled. Define
$\mathcal{D}^{\mathrm{rd}}=\{\mathcal{G}_{1}^{\mathrm{rd}},\ldots,\mathcal{G}_{p}^{\mathrm{rd}}\},\qquad\mathcal{G}_{i}^{\mathrm{rd}}=\{\mathbf{R}_{i}^{k}\mathbf{x}_{i}^{k}+\mathbf{b}_{i}^{k}\}_{k=1}^{\ell}.$
Here, the superscript “rd” denotes randomly transformed representation.
Notably, Assumptions 1 and 2 still hold for $\mathcal{D}^{\mathrm{rd}}$. Moreover, in most cases, data points generated by randomly sampled transformations are considered unseen data with respect to the original dataset, which we regard as seen data. Based on this observation, we propose leveraging the equivariance property to map both seen and unseen samples to a canonical representation, which serves as a standardized representation for each group: $\mathcal{D}^{\mathrm{cn}}=\{\mathbf{x}_{1}^{\mathrm{cn}},\mathbf{x}_{2}^{\mathrm{cn}},\cdots,\mathbf{x}_{p}^{\mathrm{cn}}\},$ where the superscript “cn” denotes canonical representation. Each $\mathbf{x}_{i}^{\mathrm{cn}}$ denotes the canonical representation shared by both $\mathcal{G}_{i}$ and its transformed counterpart $\mathcal{G}_{i}^{\mathrm{rd}}$, such that
$\forall\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists\mathbf{R},\mathbf{b},\quad\mathbf{x}=\mathbf{R}\mathbf{x}_{i}^{\mathrm{cn}}+\mathbf{b}.$
Therefore, for any $\mathrm{SE}(3)$-equivariant function $f$, it holds that:
$\forall\,\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists\,\mathbf{R},\mathbf{b},\quad f(\mathbf{x})=\mathbf{R}f(\mathbf{x}_{i}^{\mathrm{cn}})+\mathbf{b}.$ (3)
This implies that the equivariance property of $f$ allows $\mathrm{SE}(3)$ transformations to be factored out from the original dataset, which can in turn benefit the training of a robot policy $\pi$. By leveraging this property and estimating the transformation parameters $\mathbf{R}$ and $\mathbf{b}$ as described in Equation (3), the policy $\pi$ can generalize any observation within the equivariant group as
$\forall\mathbf{x}\in\mathcal{G}_{i}\cup\mathcal{G}_{i}^{\mathrm{rd}},\ \exists\mathbf{R},\mathbf{b},\quad\pi(\mathbf{x})=\mathbf{R}\pi(\mathbf{x}_{i}^{\mathrm{cn}})+\mathbf{b}.$ (4)
Consequently, once the transformation parameters are determined, both seen and unseen observations, whether from $\mathcal{D}$ or $\mathcal{D}^{\mathrm{rd}}$, can be reliably mapped to a canonical form. The policy can thus operate consistently on the canonical representation, allowing the policy to achieve generalizable performance.
III-D Differences Between Canonical Representation and SE(3) Data Augmentation
To better understand the benefits of canonical representations, we provide a conceptual comparison with $\mathrm{SE}(3)$ data augmentation, which is a commonly used technique for improving generalization to rigid pose variations in the data.
As illustrated in Fig. 2(a), $\mathrm{SE}(3)$ data augmentation operates by applying random rigid transformations to the input data. The augmented samples, together with the original data, are then used to extract observation features, with the aim of improving inference-time robustness by exposing the model to diverse poses. In contrast, Fig. 2(b) shows that canonical representation explicitly normalizes inputs to a consistent form, regardless of their original transformation, before passing them to the downstream network.
However, data augmentation has fundamental limitations in pose generalization. Since the augmented samples are finite and sampled from a subset of all possible rigid transformations, the trained policy may fail to generalize to unseen poses that were not covered during training (see the ’Rollout’ stage in Fig. 2(a)). This leads to degraded performance when encountering new viewpoints or spatial configurations of the same object. In contrast, the canonical representation approach addresses this limitation by utilizing an equivariant mapping that transforms any instance in the same equivariant group into a unique canonical pose. As a result, the network always observes the same canonicalized input, enabling better generalization even to unseen but transformation-equivalent inputs (see the ’Rollout’ stage in Fig. 2(b)).
Formally, canonical representation can be seen as a form of pre-alignment that enforces invariance to a certain group of transformations. Instead of requiring the network to learn this invariance implicitly through data diversity, the equivariant transformation module explicitly reduces the input variation. This enables the policy to learn action mappings from a standardized input space.
To enhance readability, Table I summarizes the key notation used throughout the paper before we proceed to the Method section.
TABLE I: Key notation used in the paper.
Symbol Meaning
$\mathcal{D},\;\mathcal{G}$ A set of 3D point clouds and an equivariant group (also a subset of $\mathcal{D}$).
$\mathbf{R}\in\mathrm{SO}(3),\;\mathbf{b}\in\mathbb{R}^{3}$ Rotation matrix and translation vector.
$\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{3\times N}$ The $\ell$-th point cloud sample in group $\mathcal{G}_{i}$; a $3\times N$ array whose $N$ columns are 3D points.
$\mathbf{x},\;\mathbf{y}$ Generic point cloud samples.
$\mathbf{o},\;\mathbf{a},\;\mathbf{s}$ Observation, action, and robot proprioceptive state.
$f,\;\pi$ Equivariant function and policy mapping.
“rd”, “cn”, “mn”, “de” Randomly transformed representation, canonical representation, mean operator, and decentered representation.
“pos”, “ori”, “grip” End-effector position, orientation, and gripper open width.